Determine the complex number z which satisfies the equations |z + 3i| = |z + 5 − 2i| and |z − 4i| = |z + 2i| simultaneously
1
Expert's answer
2020-05-19T08:22:13-0400

We know that:

"|z|=\sqrt{Re^2(z)+Im^2(z)}"

So, we have:

"\begin{cases}
Re^2(z)+(Im(z)+3)^2=(Re(z)+5)^2+(Im(z)-2)^2 \\
Re^2(z)+(Im(z)-4)^2=Re^2(z)+(Im(z)+2)^2 \\
\end{cases}"

From the second equation it follows that

"Im(z)=1"

Substituting into the first equation

"Re(z)=-1=>z=-1+i"